### v1

# From Stochasticity to Signal: A Bayesian Latent State Model for Reliable Measurement with LLMs

Abstract
Large Language Models (LLMs) are increasingly used to automate classification tasks in business, such as analyzing customer satisfaction from text. However, the inherent stochasticity of LLMs, in terms of their tendency to produce different outputs for the same input, creates a significant measurement error problem that is often neglected with a single round of output, or addressed with ad-hoc methods like majority voting. Such naive approaches fail to quantify uncertainty and can produce biased estimates of population-level metrics. In this paper, we propose a principled solution by reframing LLM variability as a statistical measurement error problem and introducing a Bayesian latent state model to address it. Our model treats the true classification (e.g., customer dissatisfaction) as an unobserved latent variable and the multiple LLM ratings as noisy measurements of this state. This framework allows for the simultaneous estimation of the LLM’s false positive and false negative error rates, the underlying base rate of the phenomenon in the population, the posterior probability of the true state for each individual observation, and the causal impact of a business intervention, if any, on the latent state. Through simulation studies, we demonstrate that our model accurately recovers true parameters where naive methods fail. We conclude that this methodology provides a general and reliable framework for converting noisy, probabilistic outputs from LLMs into accurate and actionable insights for scientific and business applications.
keywords:
Large Language Models (LLMs),Measurement Error, Bayesian Latent Variable Model, Text Classification, Causal Inference, Customer Satisfaction, Noisy Labels
1 Introduction
The business imperative for scalable customer satisfaction monitoring is clear, yet manual review of qualitative data like call transcripts remains a significant bottleneck. Gathering human expert ratings is expensive and time-consuming. Large Language Models (LLMs) have emerged as a natural and powerful tool for automating this classification task, exhibiting good alignment with human judgments on certain tasks (Bavaresco et al., 2024). However, a fundamental characteristic of LLMs, their probabilistic nature, presents a critical challenge. The stochasticity, while a feature for generating diverse text, becomes a source of measurement error in classification contexts where consistency is paramount. For example, if an LLM provides five ratings for a single call transcript and returns the set $\{1,0,1,1,0\}$, the true satisfaction state is ambiguous.
Naive aggregation methods, such as using a single round of LLM output as the ground truth, or taking the majority vote or the average from multiple rounds of outputs, are insufficient. They fail to model the underlying error process, cannot quantify uncertainty, and do not easily incorporate prior knowledge or external factors.
When the objective of the classification task is an important business outcome (e.g., user satisfaction) on which the impact of a business product or feature is of interest (e.g., a customer service workflow), the inability to characterize the measurement error by the naive methods hinders the causal evaluations of high-stake decisions.
In this paper, we address these challenges by proposing a Bayesian latent state model. We reframe the problem of LLM variability as a solvable measurement error problem. Our model achieves three key goals: (1) it estimates the population’s base dissatisfaction rate and the LLM’s intrinsic error rates directly from the data; (2) it quantifies the uncertainty for each individual classification; and (3) it provides a framework for measuring the impact of business interventions on customer satisfaction. This paper proceeds as follows: Section 2 reviews related work, Section 3 details our model, Sections 4 and 5 provide validations via simulation, and Section 6 concludes with discussion.
2 Background & Related Work
2.1 The Proliferation of Large Language Models in Automated Text Classification
The advent of Large Language Models (LLMs), powered by the transformer architecture, represents a paradigm shift in natural language processing (NLP). Pre-trained on vast text corpora, these models have demonstrated a remarkable capacity to understand and generate human-like text, enabling them to perform a wide array of tasks, including text classification, with minimal task-specific training. The application of LLMs for classification typically falls into three regimes: zero-shot learning, where the model classifies text based on a prompt without examples; few-shot learning, which improves performance by including a handful of examples in the prompt; and fine-tuning, which achieves state-of-the-art results by updating model weights on a larger labeled dataset (Kostina et al., 2025). While fine-tuning is powerful, many practical applications in business and science rely on the flexibility and low deployment cost of prompting-based methods (zero-shot and few-shot learning), making the variability inherent in this approach a particularly salient issue (Lu et al., 2021; Liu et al., 2023).
In business analytics and customer intelligence, LLMs are widely used to automate the analysis of unstructured text from sources like customer reviews, support tickets, and social media (Ipa et al., 2024; Ghatora et al., 2024; Wang et al., 2024). This includes tasks such as sentiment analysis to gauge customer opinion, theme extraction to identify recurring issues like "product quality" or "customer service," and competitive intelligence to monitor brand perception. By moving beyond simple star ratings to a deeper, qualitative understanding of customer feedback, businesses can make more informed, customer-centric decisions (Bharadwaj, 2023). Similarly, in the social sciences, LLMs are emerging as a tool to accelerate traditionally labor-intensive qualitative data analysis (Bano et al., 2024). Researchers are using LLMs to automate the "coding" of interview transcripts and open-ended survey responses, which promises to increase the scale and speed of qualitative inquiry. While this automation offers benefits such as reducing researcher subjectivity, the literature also highlights critical challenges, including a heavy reliance on prompt engineering and the risk of model "hallucinations," which points toward an underlying problem of reliability and measurement (Barros et al., 2025). The widespread adoption of LLMs in these applied domains has created a quiet measurement crisis: the models are often treated as deterministic black boxes, with little attention paid to the statistical implications of their inherent variability.
2.2 A Foundational Challenge: Probabilistic Outputs and Inherent Stochasticity as Measurement Error
At their core, LLMs are probabilistic models that learn a distribution over a vocabulary of tokens (Wang et al., 2023). The variability in their output stems from the decoding strategy applied to this distribution. Techniques such as setting the temperature parameter and top-$k$ sampling introduce randomness to generate diverse and creative text (Kostina et al., 2025). While this stochasticity is a feature for generative tasks, it becomes a significant challenge in classification contexts where consistency and accuracy are paramount (Bender et al., 2021; Wang et al., 2022). When an LLM is tasked with classifying an item that has a single, true underlying state (e.g., a customer is either dissatisfied or not), its output variability ceases to be a feature and must be treated as a source of measurement error (Atil et al., 2024; Ludwig et al., 2025; Herrera-Poyatos et al., 2025). Addressing this probabilistic nature of outputs from LLMs is analogous to resolving the conflicts from inconsistent human ratings, where metrics and methods like inter-rater reliability, truth inference, and multi-annotator competence estimation were developed to infer the most likely true label (Hallgren, 2012; Li et al., 2019; Martín-Morató and Mesaros, 2023).
This perspective also directly connects the problem of LLM variability to the well-established machine learning field of learning with noisy labels (LNL). When an LLM is used to annotate a dataset, its stochastic nature makes it a noisy labeler (Baumann et al., 2025; Lopez et al., 2025). Recent work has demonstrated that this LLM-generated label noise can introduce systematic, prevalence-dependent bias into the evaluation of downstream models, underscoring the severity of the problem (Baumann et al., 2025; Voutsa et al., 2025). This paper builds on the insight that instead of attempting to eliminate stochasticity (e.g., by setting the temperature to zero, which discards uncertainty information), one can leverage it. By collecting multiple ratings for the same item, we obtain a richer signal that contains information about the LLM’s classification uncertainty for that specific item. This transforms the problem from viewing stochasticity as a bug to be fixed into treating it as a signal to be modeled within a principled statistical framework.
2.3 Bias Correction using Econometric Frameworks
To rigorously address the measurement error and systematic bias inherent in Large Language Model (LLM) outputs when used in downstream estimation, recent research has proposed an applied econometric framework (Ludwig et al., 2025). This framework formally treats LLM-generated labels not as ground truth, but as imperfect, noisy surrogates for the desired, expensive-to-collect true measurements, a problem long studied in econometrics.
The constructive solution offered by this econometric approach is the use of a small "gold-standard" validation sample where the ground truth is measured directly. This validation data allows for the empirical modeling of the LLM’s error structure, such as its misclassification probabilities. By integrating this error model, researchers can construct bias-corrected estimators for target parameters. This procedure restores desirable asymptotic statistical properties like consistency and asymptotic normality, which are lost when using naive "plug-in" estimates that treat LLM outputs as perfect measurements. Furthermore, under certain conditions, incorporating the imperfect LLM outputs alongside the validation data can improve the precision of estimates by leveraging the vast information encoded in the LLM’s pre-training to amplify the signal from the validation sample.
However, this framework’s effectiveness is contingent on the ability to obtain a validation sample that is sufficiently sized, representative, and trustworthy to serve as a reliable gold standard for calibrating the LLM’s performance in the specific task of interest. In some business context, it might be difficult or even prohibited to obtain such high-quality validation sample (Sheng et al., 2008; Roh et al., 2019; Sambasivan et al., 2021).
2.4 Theoretical Grounding in Learning with Noisy Labels (LNL)
The LNL literature provides a formal language for understanding and modeling the measurement error of LLMs. The central challenge in LNL is training a robust classifier on a dataset where labels are corrupted, a common issue in large-scale data collection. A key difficulty is that deep neural networks have a high capacity to overfit or "memorize" these noisy labels, which degrades their ability to generalize to new data (Karimi et al., 2020).
A cornerstone of LNL is the noise transition matrix, $T$, where each element $T_{ij}=P(\tilde{y}=j|y=i)$ represents the probability that a true label $i$ is observed as a noisy label $j$ (Zhang et al., 2021). This formalism directly maps to the parameters in our proposed model for a binary classification task. The LLM’s false positive rate, $\epsilon_{0}$, is the probability of it assigning a "dissatisfied" label (e.g., 1) when the true state is "satisfied" (0), or $T_{01}$. Similarly, the false negative rate, $\epsilon_{1}$, is the probability of it assigning a "satisfied" label (0) when the true state is "dissatisfied" (1), or $T_{10}$. Our base model operates under the common assumption of Class-Conditional Noise (CCN), where the error probability depends only on the true class, not the specific features of the instance. We also propose an extended model to relax this assumption to explain the correlation between the error probabilities for the same instance using a ‘difficulty’ label.
While the LNL literature provides the formal tools for modeling such error processes, its primary goal is typically to use this model to train a separate, robust downstream classifier (Kim et al., 2024). In contrast, our work uses these concepts for a different inferential goal: the measurement and characterization of the LLM’s performance and the true underlying states themselves. The estimation of the error rates and true labels is the end goal, not a means to train another model.
2.5 A Bayesian extension of a Latent Class Model for Diagnostic Accuracy and Causal Inference
For decades, Latent Class Models (LCAs) have been the standard methodology in epidemiology and biostatistics for a critical and analogous problem: evaluating the accuracy of one or more diagnostic tests when no perfect "gold standard" reference test is available (Cheung et al., 2021). In this context, the true disease status of a patient is treated as the unobserved latent class. The model uses the patterns of agreement and disagreement among multiple imperfect tests across a population to simultaneously estimate the prevalence of the disease and the accuracy of each test.
The problem of evaluating an LLM’s classification performance without a perfectly labeled dataset is structurally identical to this medical diagnostics problem. A diagnostic test takes a patient and produces a noisy binary signal. The LLM takes a call transcript and produces a noisy binary signal as well. In both scenarios, the true underlying state is unknown, and the statistical task is to use the patterns of these noisy signals to infer the properties of both the population (base rate) and the signal generator (the test or the LLM). The Dawid-Skene model, a specific instance of the LCA family, was originally developed to tackle this task (Dawid and Skene, 1979). Our proposed method is a direct Bayesian generalized linear model (GLM) version of the Dawid-Skene model, allowing the flexible incorporation of additional covariates to help improve the accuracy of the latent state. When an intervention is a part of the attributes for each patient or caller, our method also enables the causal inference of its impact when the intervention is randomized or observed along with an admissible set of confounders (Pearl, 2010). The Bayesian nature of the method further extends its ability to include prior knowledge and quantify uncertainty.
2.6 Synthesis and Contribution: A New Framework for Reliable Measurement with LLMs
This paper addresses a critical gap at the intersection of machine learning and applied science. While LLMs are widely deployed for classification, their inherent stochasticity creates a fundamental measurement problem that is often ignored or addressed with ad-hoc methods like majority voting, which fail to quantify uncertainty or correct for systematic bias (Bender et al., 2021; Wang et al., 2022). Although the LNL literature offers tools for modeling error processes (Zhang et al., 2021) and the latent variable modeling literature provides a framework for inference without a gold standard (Pepe and Janes, 2007), no work has synthesized these traditions to specifically address the measurement error arising from the inherent stochasticity of modern generative AI, especially in the absence of sufficiently sized, representative, and trustworthy validation data.
Our contribution is a principled, statistically rigorous framework for converting the noisy, probabilistic outputs of LLMs into reliable and actionable insights. The repeated sampling from LLM outputs replaces the need for the often unattainable gold standard data especially when dealing with huge amounts of customer interactions. By treating the true classification as a latent variable and the multiple LLM ratings as noisy indicators, our Bayesian model offers several key advantages over naive approaches. It simultaneously estimates the LLM’s own false positive and false negative error rates alongside the population base rate of the phenomenon of interest. It provides principled uncertainty quantification through full posterior distributions for every parameter and for the latent state of each individual observation. Finally, the model’s parameters are directly interpretable and the framework is easily extended to include covariates and measure the effects of business interventions, making it a powerful tool for both measurement and causal inference. We propose a new paradigm for using LLMs as measurement tools: one that embraces their probabilistic nature and subjects their outputs to the same level of statistical rigor expected of any scientific instrument.
3 Methodology: A Bayesian Latent State Model
3.1 Model Overview & Goals
Our model is designed to estimate the underlying true satisfaction states from noisy LLM ratings. The primary objectives are to estimate the base rates of dissatisfaction, the AI’s error rates, the posterior probability of dissatisfaction for each individual call, and to measure the effect of treatments or interventions.
3.2 Model Specification
Let $c=1,\dots,C$ index the calls. For each call, we define the following:
  • Latent State: $D_{c}\in\{0,1\}$ is the true, unobserved dissatisfaction state for call $c$, where $D_{c}=1$ indicates dissatisfaction.
  • Observed Data: $R_{c,i}$ is the $i$-th rating from the LLM for call $c$, where $i=1,\dots,N_{c}$. Let $k_{c}=\sum_{i=1}^{N_{c}}R_{c,i}$ be the total number of "dissatisfied" ratings for call $c$.
  • Covariates (Optional): $X_{c}$ is a set of attributes for call $c$ or its caller that helps predict the latent class, or a set of admissible confounders for inferring the effect of a business intervention.
  • Treatment (Optional): $T_{c}\in\{0,1\}$ is the business intervention for the caller of $c$ (e.g. two variants of customer support programs), of which the causal impact is of interest. We assume the treatment is randomized in an experiment, or observed along with an admissible set of confounders $X_{c}$.
The model estimates the following parameters:
  • $\theta_{c}=P(D_{c}=1)$, the base rate of dissatisfaction. We model this using a logit link to incorporate an intercept ($\theta$), covariates ($X_{c}$), and a treatment effect ($\tau$):
    $\text{logit}(\theta_{c})=\theta+X_{c}\beta+\tau T_{c}$ (1)
  • $\epsilon_{0}=P(R_{c,i}=1|D_{c}=0)$, the LLM’s False Positive Rate.
  • $\epsilon_{1}=P(R_{c,i}=0|D_{c}=1)$, the LLM’s False Negative Rate.
  • $\eta=E[D_{c}(T_{c}=1)-D_{c}(T_{c}=0)]$, the average treatment effect (ATE) of the business intervention on an additive scale.
Note: we assume constant false positive rate and false negative rate for simplicity, but it is also possible to relax this constraint by imposing models on the two errors rates to be varied by each single call. This extension, discussed in Section 3.6, helps characterize the inherent heterogeneous difficulty in classifying each call, and the correlated structure between ratings from the same call.
3.3 The Likelihood Function
The model for the observed data $k_{c}$ is a mixture model, conditional on the latent state $D_{c}$. The likelihood for a single call $c$ is given by:
$\begin{split}P(k_{c}\mid\theta_{c},\epsilon_{0},\epsilon_{1})={}&(1-\theta_{c})P(k_{c}\mid D_{c}=0,\epsilon_{0})\\ &+\theta_{c}P(k_{c}\mid D_{c}=1,\epsilon_{1})\end{split}$ (2)
where the conditional probabilities are binomial:
  • If the call is truly satisfactory ($D_{c}=0$), the number of positive ratings $k_{c}$ follows a binomial distribution based on the false positive rate:
    $P(k_{c}|D_{c}=0,\epsilon_{0})=\binom{N_{c}}{k_{c}}\epsilon_{0}^{k_{c}}(1-\epsilon_{0})^{N_{c}-k_{c}}$ (3)
  • If the call is truly dissatisfactory ($D_{c}=1$), the number of positive ratings $k_{c}$ follows a binomial distribution based on one minus the false negative rate:
    $P(k_{c}|D_{c}=1,\epsilon_{1})=\binom{N_{c}}{k_{c}}(1-\epsilon_{1})^{k_{c}}\epsilon_{1}^{N_{c}-k_{c}}$ (4)
3.4 Priors and Computation
We adopt a fully Bayesian approach, specifying priors for all unknown parameters:
$\theta$ $\sim\mathcal{N}(\mu_{\theta},\sigma_{\theta}^{2})$
$\beta$ $\sim\mathcal{N}(\mu_{\beta},\sigma_{\beta}^{2})$
$\tau$ $\sim\mathcal{N}(\mu_{\tau},\sigma_{\tau}^{2})$
(Weakly informative priors) (5)
$\epsilon_{0},\epsilon_{1}$ $\sim\text{Beta}(1,1)$
(Uniform, uninformative priors) (6)
The error rates are further truncated to be smaller than $0.5$ to resolve ‘label switching‘, a well-known issue with two-class mixture models (Stephens, 2000). The model is estimated using Markov Chain Monte Carlo (MCMC) methods, implemented in Stan.
3.5 Deriving Key Quantities of Interest
The posterior distributions for the primary parameters ($\theta,\beta,\tau,\epsilon_{0},\epsilon_{1}$) are obtained directly from the MCMC output. The posterior probability of dissatisfaction for each individual call $c$ is calculated using Bayes’ rule, leveraging the estimated parameters:
$\begin{split}&P(D_{c}=1|k_{c},\hat{\theta}_{c},\hat{\epsilon}_{0},\hat{\epsilon}_{1})=\\ &\quad\frac{P(k_{c}|D_{c}=1,\hat{\epsilon}_{1})\hat{\theta}_{c}}{P(k_{c}|\hat{\theta}_{c},\hat{\epsilon}_{0},\hat{\epsilon}_{1})}\end{split}$ (7)
3.6 Model Extension
One of the key model assumptions is the conditional independence assumption that the observed LLM ratings are independent given the true latent model states. Its data likelihood is governed by the LLM’s false positive rate and false negative rate, descriptive of the average model performance in categorizing calls from both latent states.
In some cases, the difficulty in categorizing different calls may be more variant. An extension to both characterize the heterogeneity in the error rates among calls, and explain the correlation between LLM ratings from the same call, is desired.
Therefore, we impose models for both error rates, $\epsilon_{0c}$ and $\epsilon_{1c}$, to depend on an observable attribute of ‘difficulty’, $H_{c}$, in classifying each call. In practice, researchers may prompt the LLM to report the difficulty level prior to the classification task. We could assume that the LLM’s difficulty rating is a noisy measurement of the true difficulty for each call, centered around its true value. Under this assumption, we could ask the LLM to output the difficulty rating in each round, and take the average to obtain an unbiased measurement of the true difficulty.
The error rates models are given by:
  • $\text{logit}(2*\epsilon_{0c})=\alpha_{0}+\gamma_{0}H_{c}$, the LLM’s False Positive Rate for each call $c$.
  • $\text{logit}(2*\epsilon_{1c})=\alpha_{1}+\gamma_{1}H_{c}$, the LLM’s False Negative Rate for each call $c$.
Note that doubling the error rates inside the logit function assumes that both error rates are smaller than 50%, i.e. the LLM outperforms a random guess in the worst case scenario. This helps prevent the ‘label switching’ issue similar to the truncation method in our base model.
The likelihood is simply updated by replacing the two error rates with the inverse logit functions of the linear models for each call. For choices of priors, we additionally assume:
$\alpha_{0}$ $\sim\mathcal{N}(\mu_{\alpha_{0}},\sigma_{\alpha_{0}}^{2})$
$\alpha_{1}$ $\sim\mathcal{N}(\mu_{\alpha_{1}},\sigma_{\alpha_{1}}^{2})$
$\gamma_{0}$ $\sim\text{LogNormal}(\mu_{\gamma_{0}},\sigma_{\gamma_{0}}^{2})$ (8)
$\gamma_{1}$ $\sim\text{LogNormal}(\mu_{\gamma_{1}},\sigma_{\gamma_{1}}^{2})$ (9)
The slopes $\gamma_{0}$ and $\gamma_{1}$ are assumed to follow lognormal distributions to indicate positive correlation between difficulty and error rates.
4 Simulation Study 1
4.1 Objective
The objective is to demonstrate that the base model can successfully recover known, true parameters from simulated data and to compare its performance against naive aggregation methods.
4.2 Simulation Design
We simulate a dataset based on the generative process assumed by our model, which includes call-level covariates and a treatment effect. The data generating process (DGP) follows these steps:
  1. 1.
    Simulate Call Characteristics: We first simulate $C=1,000$ unique calls. For each call $c$, we generate two covariates: a continuous variable $x_{c,1}\sim\mathcal{N}(0,1)$ and a binary variable $x_{c,2}\sim\text{Bernoulli}(0.4)$. Each call is also randomly assigned to a treatment group $T_{c}\sim\text{Bernoulli}(0.5)$.
  2. 2.
    Simulate Latent States: The true, latent dissatisfaction state $D_{c}$ for each call is drawn from a Bernoulli distribution, where the probability of dissatisfaction $\theta_{c}$ is determined by a logistic regression model:
    $\text{logit}(\theta_{c})=\theta+\beta_{1}x_{c,1}+\beta_{2}x_{c,2}+\tau T_{c}$
    where $\theta_{c}=P(D_{c}=1|x_{c,1},x_{c,2},T_{c})$. The true parameter values (or "ground truth") are set as follows:
    • Intercept ($\theta$): $\theta=-1.5$
    • Covariate Effects ($\bm{\beta}$): $\beta_{1}=0.75$, $\beta_{2}=-0.5$
    • Treatment Effect ($\tau$): $\tau=-0.8$
    The latent state for each call is then simulated from its unique probability: $D_{c}\sim\text{Bernoulli}(\theta_{c})$.
  3. 3.
    Simulate Observed Ratings: For each call $c$, we simulate $N_{c}=5$ independent AI ratings, denoted $R_{c,i}$ for $i\in\{1,\dots,5\}$. The probability of a "dissatisfied" rating ($R_{c,i}=1$) is conditional on the true latent state $D_{c}$ and the AI’s error characteristics. We define the false positive rate (FPR) as $\epsilon_{0}=P(R_{c,i}=1|D_{c}=0)$ and the false negative rate (FNR) as $\epsilon_{1}=P(R_{c,i}=0|D_{c}=1)$.
    The observed ratings are thus drawn from:
    $R_{c,i}|D_{c}\sim\begin{cases}\text{Bernoulli}(\epsilon_{0})&\text{if }D_{c}=0\\ \text{Bernoulli}(1-\epsilon_{1})&\text{if }D_{c}=1\end{cases}$
    For this study, we set the true error rates to $\epsilon_{0}=0.15$ and $\epsilon_{1}=0.10$.
  4. 4.
    Fit the proposed model: We fit our proposed method to the simulated dataset with $1,000\times 5=5,000$ total ratings $R_{c,i}$, along with the observed covariates and the treatment $x_{c,1},x_{c,2},T_{c}$. The prior parameters are set as follows:
    • Logistic regression: $\mu_{\theta}=-3$, $\sigma_{\theta}=2$, $\mu_{\beta}=0$, $\sigma_{\beta}=1$, $\mu_{\tau}=0.05$, $\sigma_{\tau}=0.5$
    • FPR & FNR: $\text{Beta}(1,1)$ truncated within $(0,0.5)$
  5. 5.
    Fit naive approaches for comparison: We compare the performance of our proposed method with two other naive methods:
    • Single-round AI rating: We fit the same Bayesian latent model to only the first rating of each call, $R_{c,1}$, mimicking the real-world practice of inferring the latent class based on a single-round AI rating.
    • Majority vote: We take the majority vote out of the five ratings for each call as the estimated latent class. The error rates are estimated through taking the majority vote as the true latent class. We fit to the majority vote class a frequentist logistic regression model for estimating the impact parameters.
4.3 Evaluation & Results
We evaluate the model on three primary criteria: its ability to recover the true parameters of the simulation, the accuracy of its predictions for the latent states, and its performance relative to common naive methods. Our proposed model, which constrains the error rates $\epsilon_{0},\epsilon_{1}<0.5$ to prevent label switching, successfully recovers the ground truth parameters. An initial unconstrained model demonstrated a bimodal posterior distribution, a classic sign of label switching, failing to identify the true parameters. All subsequent results refer to the constrained model.
4.3.1 Parameter Recovery
The model demonstrates a strong ability to recover the known, true parameters from the simulated data. The posterior distributions for all key parameters are centered directly around their true values, indicating that the model is well-specified and provides accurate estimates.
AI Error Rates ($\bm{\epsilon_{0},\epsilon_{1}}$): The model accurately estimates the AI’s reliability. The posterior mean for the False Positive Rate ($\epsilon_{0}$) was 0.159, closely matching the true value of 0.15. Similarly, the posterior mean for the False Negative Rate ($\epsilon_{1}$) was 0.105, nearly identical to the true value of 0.10. Figure 1 visually confirms that the posterior distributions are concentrated around the ground truth values.
Treatment Effects ($\bm{\tau,\eta}$): The model also successfully recovers the causal parameters. The posterior mean for the treatment effect on the log-odds scale ($\tau$) was -0.711, closely tracking the true value of -0.8. For the Average Treatment Effect (ATE) on the probability scale ($\eta$), the posterior mean was -0.0764, which is very close to the true ATE of -0.0871. Figure 2 shows the posterior distributions for these effects are correctly centered.
Figure 1: Posterior distributions for the AI’s False Positive Rate ($\epsilon_{0}$) and False Negative Rate ($\epsilon_{1}$). The red dashed lines indicate the true values used in the simulation. The model’s posteriors are correctly centered on these true values.
Figure 2: Posterior distributions for the treatment effect coefficient ($\tau$) and the Average Treatment Effect ($\eta$). The red dashed lines indicate the true values. The model successfully recovers the ground truth for both parameters.
4.3.2 Accuracy of Latent State Estimation
Beyond recovering aggregate parameters, the model excels at estimating the true, unobserved dissatisfaction state for each individual call. The mean posterior probability of dissatisfaction, $P(D_{c}=1)$, shows a very high correlation of 0.959 with the true latent state $D_{c}$. Further, using these posterior probabilities to classify calls yields an Area Under the Curve (AUC) of 0.9985, indicating near-perfect classification performance. Figure 3 illustrates this clear separation, where calls with a true state of 1 have posterior probabilities clustering near 1, and vice-versa for calls with a true state of 0.
Figure 3: Mean posterior probability of dissatisfaction for each call, grouped by the true latent state. The model’s estimates clearly distinguish between the two true states.
4.3.3 Comparison with Naive Methods
We compared our proposed model, which leverages $N=5$ ratings per call, against two common, simpler approaches: (1) a "Naive N=1" approach that applies the same model to only the first AI rating, simulating the common practice of using a single LLM output, and (2) a "Majority Vote" approach that classifies a call as dissatisfied if 3 or more of the 5 ratings are positive, and computes the error rates and impact parameters taking the majority vote as if it is the true latent state.
As shown in Table 1, the proposed model with five LLM ratings for each call significantly outperforms fitting the same model with one single rating for each call. The Naive N=1 approach yields estimates for the error rates and treatment effects that are substantially biased. For instance, it overestimates the FNR by more than double (0.235 vs. a true value of 0.1) and underestimates the magnitude of the treatment effect $\tau$ by over 30% (-0.540 vs. -0.8). Its ability to recover the true latent state is also poor, with a correlation of only 0.575.
The Majority Vote approach performs better, achieving a high correlation of 0.915 with the true state, which is intuitive given that the true FPR and FNR are quite low. However, our proposed model achieves an even higher correlation (0.959), demonstrating its superior ability to synthesize the information contained in the multiple ratings into a more accurate estimate of the latent state. Beyond that, both methods also accurately recover important parameters like the error rates and impact parameters. However, our Bayesian approach also quantifies uncertainty within these parameters through posterior probabilities and credible intervals, but the majority vote only provides a single point estimate. Empirical researchers need to perform bootstrap sampling, or generate multiple LLM output datasets, for calculating the confidence intervals for the parameters recovered by the majority vote approach, especially for the impact parameters estimated in the second stage of a logistic regression, where the uncertainty from the first voting stage is not accounted for in the regression model. This highlights the value of explicitly modeling the measurement error process rather than relying on simple aggregation.
Table 1: Model comparison showing parameter recovery and accuracy.
Metric True Proposed Naive Majority
Value (N=5) (N=1) Vote
Corr. $D_{\text{true}}$ - 0.959 0.575 0.915
FPR 0.15 0.159 0.185 0.148
FNR 0.10 0.105 0.235 0.137
$\tau$ -0.80 -0.711 -0.540 -0.686
$\eta$ -0.087 -0.076 -0.053 -0.084
5 Simulation Study 2
5.1 Objective
The objective is to demonstrate the superior performance of the extended model when the difficulty level in the classification task is highly variant among the collected calls. It also discusses how the model accuracy evolves as the number of LLM rounds of ratings increases.
5.2 Simulation Design
We simulate a dataset similar to the previous design in Section 4.2 with the following updates:
  1. 1.
    Simulate Call Characteristics: We simulate $C=1,500$ unique calls.
  2. 2.
    Simulate Observed Difficulty: The observed difficulty $H_{c}$ for each call is drawn from a uniform distribution between 0 and 1.
  3. 3.
    Simulate Observed Ratings: For each call $c$, we simulate $N_{c}=10$ independent AI ratings, denoted $R_{c,i}$ for $i\in\{1,\dots,10\}$. The probability of a "dissatisfied" rating ($R_{c,i}=1$) is conditional on the true latent state $D_{c}$ and the observed difficulty $H_{c}$. We define the false positive rate (FPR) as $\epsilon_{0c}=P(R_{c,i}=1|D_{c}=0,H_{c})$ and the false negative rate (FNR) as $\epsilon_{1c}=P(R_{c,i}=0|D_{c}=1,H_{c})$. They are determined by the logistic regression models:
    • $\text{logit}(2*\epsilon_{0c})=\alpha_{0}+\gamma_{0}H_{c}$, the LLM’s False Positive Rate for each call $c$.
    • $\text{logit}(2*\epsilon_{1c})=\alpha_{1}+\gamma_{1}H_{c}$, the LLM’s False Negative Rate for each call $c$.
    The true parameter values are set as follows to allow the error rates to be varied by the difficulty levels:
    • Intercepts: $\alpha_{0}=\alpha_{1}=-1.5$
    • Slopes: $\gamma_{0}=3.5$, $\gamma_{1}=4.0$
    The observed ratings are thus drawn from:
    $R_{c,i}|D_{c}\sim\begin{cases}\text{Bernoulli}(\epsilon_{0c})&\text{if }D_{c}=0\\ \text{Bernoulli}(1-\epsilon_{1c})&\text{if }D_{c}=1\end{cases}$
  4. 4.
    Fit the proposed model: The prior parameters for the logistic regressions of the error rates are set as follows:
    • Intercepts: $\mu_{\alpha_{0}}=-1.5$, $\sigma_{\alpha_{0}}=1$, $\mu_{\alpha_{1}}=-1.5$, $\sigma_{\alpha_{1}}=1$
    • Slopes: $\mu_{\gamma_{0}}=0$, $\sigma_{\gamma_{0}}=0.5$, $\mu_{\gamma_{1}}=0$, $\sigma_{\gamma_{1}}=0.5$
  5. 5.
    Fit other approaches for comparison: We compare the performance of our extended methods, our base methods, and the naive methods, varying the number of LLM ratings as input:
    • Heterogeneous Measurement Error model (Extension method): We fit the extended model to the observed datasets taking all 10 ratings, first 5 ratings, or only the first rating of each call.
    • Homogeneous Measurement Error model (Base method): We fit the base model to the observed datasets taking all 10 ratings, first 5 ratings, or only the first rating of each call.
    • Majority vote: We take the majority vote out of the 10 ratings for each call as the estimated latent class, without fitting the proposed models.
5.3 Model Comparison
We use ’heterogeneous’ or ’homogeneous’ to denote the extended method or the base method that models the error rates heterogeneously or homogeneously for calls.
Shown in Table 2, the extended model with ten LLM ratings for each call outperforms the rest of the methods in terms of the correlation with the latent dissatisfaction status (0.803). When the heterogeneity in error rates is captured, the extended model with five rounds of LLM outputs outperforms the base model with ten rounds of outputs. Both extended and base methods have improved accuracies in estimating the true latent dissatisfaction as the input rounds increases. The majority vote approach with all ten rounds of input merely outperforms the proposed methods with one single round of output.
When the impact parameters are of interest, Table 3 shows that the performance of the majority vote approach further deteriorates by providing the least accurate estimates. The extended method produces more accurate estimates than the base method conditioning on all three choices of output rounds. However, the accuracy in recovering the impact parameters does not increase along with the output rounds. The extended method with five rounds of LLM output provides the best estimate of both impact parameters across all tested model settings.
This phenomenon illustrates a critical tradeoff in mixture models with dual purposes of inference. The likelihood is a mixture of two components: a latent outcome regression component (the logistic regression modeling the true state, where $\tau$ and $\eta$ are estimated) and a measurement error component (the binomial likelihood of the observed ratings, conditional on the true state). As the number of LLM output rounds increases, the binomial likelihood becomes substantially more concentrated. Consequently, the model’s objective becomes dominated by fitting the highly flexible individualized error rate functions to the precise $k_{c}/N_{c}$ ratios observed for each unit. This can lead to a form of parameter confounding, where variance that should be attributed to the treatment effect is instead absorbed by the measurement component. The model mis-attributes a portion of the true causal effect to nuanced patterns in measurement error, resulting in biased estimates for $\tau$ and $\eta$. The N=5 scenario appears to strike an optimal balance, where the measurement component is informative but does not completely overwhelm the latent outcome regression component of the likelihood.
Table 2: Model comparison showing correlation with latent states.
Method Corr. with $D_{\text{true}}$
Heterogeneous (N=10) 0.803
Homogeneous (N=10) 0.702
Heterogeneous (N=5) 0.723
Homogeneous (N=5) 0.670
Heterogeneous (N=1) 0.485
Homogeneous (N=1) 0.455
Majority Vote (N=10) 0.662
Table 3: Model comparison showing impact parameter recoveries.
Method $\tau$ $\eta$
True Value -0.80 -0.086
Heterogeneous (N=10) -0.51 -0.059
Homogeneous (N=10) -0.31 -0.054
Heterogeneous (N=5) -0.69 -0.072
Homogeneous (N=5) -0.48 -0.070
Heterogeneous (N=1) -0.52 -0.050
Homogeneous (N=1) -0.43 -0.049
Majority Vote (N = 10) -0.29 -0.041
6 Discussion and Conclusion
6.1 Summary of Contributions
We address the problem of measurement error inherent in probabilistic LLMs by proposing a principled Bayesian latent state model. Our approach successfully estimates three sets of underlying parameters of interest while rigorously quantifying uncertainty: the outcome target of the classification task (e.g. satisfaction), the AI measurement error, and the impact of a randomized business intervention on the outcome.
With the recovery of the three sets of parameters, our model serves as a valuable tool for diverse purposes. It improves validity in measuring the latent customer outcomes represented by transcripts or texts, especially when a gold-standard validation set is unavailable in zero-shot and few-shot learning. It outputs the measurement errors by the LLM as valuable information to describe or compare the performances of LLM across scenarios of testing interest (e.g., varied by prompts, tasks, temperatures, or models). It can be integrated in the analytical workflow in A/B tests that infers the impact of business products or features on the latent customer outcomes.
Our model demonstrates its significant improvements over other naive approaches, delivering two key messages for empirical researchers and analysts: (1) running multiple rounds of LLM for the same transcripts provides the data input required to address measurement error; (2) taking a modeling approach for the measurement error process effectively helps with accurate recovery of the parameters of interest with adequate uncertainty quantification.
The extended model further tackles the potential heterogeneity of the difficulty levels in the classification task among calls. It suggests a practical resolution for empirical researchers that incorporates into the prompt a side task of having LLMs rate the difficulty level along with the classification. Processing this additional information, the model preserves its accuracy with more complex data. The simulated comparison between models under different numbers of LLM output rounds highlights the trade-off between the recoveries of different sets of parameters. When recovering the true state of satisfaction to the maximum accuracy is the paramount purpose, it is better to run more rounds of the LLM within computational capacity; when the impact parameters are of equal interest, an adequate ‘medium’ number of LLM rounds is preferred (e.g., N = 5 instead of N = 1 or 10 in Simulation Study 2).
6.2 Future Work & Implications
Future extensions could adapt the model for multinomial or ordinal ratings (e.g., 1-5 stars). As probabilistic AI becomes a standard tool for measurement, rigorous statistical methods are essential for turning noisy outputs of diverse complex data-generating structures into reliable, actionable insights.
While the proposed model was compared to some prevalent naive aggregation methods to demonstrate its superior performance, it is also possible to conduct a more comprehensive investigation of our model performances combined with, or in contrast to, different state-of-the-art aggregation approaches motivated by resolving inconsistent human expert ratings (e.g., multi-annotator competence estimation). For another example, when multiple LLM runs produce conflicting ratings, a ’meta-LLM’ or a follow-up prompt could be tasked with adjudicating the conflict, similar to how human raters resolve disagreements.
Finally, our framework enables the evaluation and comparison of LLM performances through the error rates across different settings. Future applications of our proposed methods could involve real case studies comparing the error rates varied by different types of tasks, prompts, temperature settings, etc.

### v3

@@
 We use ’heterogeneous’ or ’homogeneous’ to denote the extended method or the base method that models the error rates heterogeneously or homogeneously for calls.
-Shown in Table 2, the extended model with ten LLM ratings for each call outperforms the rest of the methods in terms of the correlation with the latent dissatisfaction status (0.803).
-When the heterogeneity in error rates is captured, the extended model with five rounds of LLM outputs outperforms the base model with ten rounds of outputs.
-Both extended and base methods have improved accuracies in estimating the true latent dissatisfaction as the input rounds increases.
-The majority vote approach with all ten rounds of input merely outperforms the proposed methods with one single round of output.
-When the impact parameters are of interest, Table 3 shows that the performance of the majority vote approach further deteriorates by providing the least accurate estimates.
-The extended method produces more accurate estimates than the base method conditioning on all three choices of output rounds.
-However, the accuracy in recovering the impact parameters does not increase along with the output rounds.
-The extended method with five rounds of LLM output provides the best estimate of both impact parameters across all tested model settings.
+As shown in Table 2, the extended model with ten LLM ratings for each call outperforms the rest of the methods in terms of the correlation with the latent dissatisfaction status (0.798).
+When the heterogeneity in error rates is captured, the extended model consistently outperforms the base model across different settings of input rounds.
+Both extended and base methods have improved accuracies in estimating the true latent dissatisfaction as the input rounds increase.
+The majority vote approach with all ten rounds of input exhibits a comparable performance to the proposed methods with five rounds of output.
+When the impact parameters are of interest, Table 3 shows that the performance of the majority vote approach further deteriorates, providing less accurate estimates than our proposed methods with five rounds of output.
+The extended method produces overall more accurate estimates than the base method conditioning on all three choices of output rounds.
+However, the accuracy in recovering the impact parameters does not necessarily increase along with the output rounds.
+For example, both the base method and the extended method with five rounds of LLM output outperform their counterparts with ten rounds of LLM output in terms of the recovery $\tau$.
 This phenomenon illustrates a critical tradeoff in mixture models with dual purposes of inference.
@@
 As the number of LLM output rounds increases, the binomial likelihood becomes substantially more concentrated.
-Consequently, the model’s objective becomes dominated by fitting the highly flexible individualized error rate functions to the precise $k_{c}/N_{c}$ ratios observed for each unit.
+Consequently, the model’s objective becomes more dominated by fitting the highly flexible individualized error rate functions to the precise $k_{c}/N_{c}$ ratios observed for each unit.
 This can lead to a form of parameter confounding, where variance that should be attributed to the treatment effect is instead absorbed by the measurement component.
-The model mis-attributes a portion of the true causal effect to nuanced patterns in measurement error, resulting in biased estimates for $\tau$ and $\eta$.
-The N=5 scenario appears to strike an optimal balance, where the measurement component is informative but does not completely overwhelm the latent outcome regression component of the likelihood.
+The model misattributes a portion of the true causal effect to nuanced patterns in measurement error, resulting in more biased estimates for $\tau$ and $\eta$.
+In this case, a ‘medium’ number of LLM rounds might be preferred with the dual purpose of estimating the latent states and impact parameters.
 Table 2: Model comparison showing correlation with latent states.
 Method Corr. with $D_{\text{true}}$
-Heterogeneous (N=10) 0.803
-Homogeneous (N=10) 0.702
-Heterogeneous (N=5) 0.723
-Homogeneous (N=5) 0.670
-Heterogeneous (N=1) 0.485
-Homogeneous (N=1) 0.455
-Majority Vote (N=10) 0.662
+Heterogeneous (N=10) 0.798
+Homogeneous (N=10) 0.725
+Heterogeneous (N=5) 0.699
+Homogeneous (N=5) 0.627
+Heterogeneous (N=1) 0.434
+Homogeneous (N=1) 0.394
+Majority Vote (N=10) 0.659
 Table 3: Model comparison showing impact parameter recoveries.
 Method $\tau$ $\eta$
-True Value -0.80 -0.086
-Heterogeneous (N=10) -0.51 -0.059
-Homogeneous (N=10) -0.31 -0.054
-Heterogeneous (N=5) -0.69 -0.072
-Homogeneous (N=5) -0.48 -0.070
-Heterogeneous (N=1) -0.52 -0.050
-Homogeneous (N=1) -0.43 -0.049
-Majority Vote (N = 10) -0.29 -0.041
+True Value -0.80 -0.085
+Heterogeneous (N=10) -0.86 -0.087
+Homogeneous (N=10) -0.61 -0.092
+Heterogeneous (N=5) -0.84 -0.077
+Homogeneous (N=5) -0.64 -0.086
+Heterogeneous (N=1) -0.71 -0.051
+Homogeneous (N=1) -0.48 -0.035
+Majority Vote (N = 10) -0.52 -0.065
 6 Discussion and Conclusion